OpenStudy (anonymous):

Let f(x) = 3x2 + x – 3 and g(x) = x2 – 5x + 1. Find f(x) – g(x

OpenStudy (anonymous):

use "^" to show exponents

OpenStudy (cggurumanjunath):

3x^2+x-3-x^2+5x-1=2x^2+6x-4

OpenStudy (anonymous):

thanks guys .

OpenStudy (cggurumanjunath):

you're welcome.

OpenStudy (anonymous):

\[3x^2+x-3-(x^2-5x+1)\] \[=3x^2+x-3-x^2+5x-1\] \[=2x^2+6x-4\]